# The modelling of the action potentials in myelinated nerve fibres

### v2

@@ 2.3 Solutions of Lieberstein model
reducing, therefore, the partial differential equation (PDE) to an ordinary differential equation (ODE) and then to use standard ODE solvers for integration in time. For integration in time, the model equations are rewritten as a system of first-order ODE’s and a standard numerical integrator is applied. In the numerical examples given in the present paper the ODEPACK FORTRAN code (see [43]) ODE solver is used through its NumPy implementation. Handling of the data and initialization of the variables is done in Python by making use of the package SciPy (see [44]) and the numerical results are analysed and visualised in the Matlab environment.
The following parameters are used: (number of spatial nodes), (time in [ms]), µ (membrane capacitance), µ (axon radius), (axoplasm resistance) while the HH equations parameters are the same as taken in [1]: and while (initial values for parameters in HH equations at ). We take in the following example the axoplasm capacitance as µ and (resistance of an axon per unit length). We take (inductivity) for the numerical example (resulting in 1 propagation velocity for the AP signal). We remark that, as noted earlier by Kaplan and Trujillo [28], have calculated axon inductance from the movement of ions as if AP velocity is at µ. However, as in our case the reference axon radius is taken smaller it makes sense to reduce the value. Initially system is taken at rest and at we generate a narrow bell-shaped pulse (“spark") for the in the middle of the space domain which generates the propagating AP. In spatial units, the length of the computation node () is 61 while the width of the computational domain in the space (from to ) is 4. For the sake of readability, the additional equations and values of physical quantities used in numerical simulation
are collected in Table 1 following [1] and [23]:
| Table 1: Parameters for finding example solutions for (8) and (9). | ||
@@
It should be stressed, that the chosen parameters do not represent any particular model nerve as these have been collected from studies describing different experimental setups or even taken as a rough estimate (like specific inductivity L). Most of the parameters are taken from [1] corresponding to giant squid axon at about while inductivity is chosen so that the AP signal would have propagation velocity of around 1 if the axon radius is 1 [µ. It can be noted that the used value of inductance is large (40 [mHcm]) compared to, for example, a solid copper wire which would have inductance of roughly 18 if the radius of the wire is 1 µ and the length is 1 . Although, clearly, the structure of the axon is very different from a solid metal fibre, the difference of roughly 6 orders of magnitude is significant. What inductivity is, in principle, is a process that resists the change of magnetic field and considering various geometrical structures (like cytoskeleton) and different physical processes interacting with each other the chosen value does not appear to be outright implausible. One can note, for example, the significant mass difference between electrons and considered ions. However, it is clear that the nature of inductivity in the context of nerve fibres needs further clarification and study. We have a process in the proposed model that behaves like inductivity, and has a dimension related to inductivity so we have to make an assumption that it is inductivity to avoid violating Occam’s razor principle – although the relatively large value of the parameter needed to explain the experimental observations hints that there might be something more going on (see also Wang et al. [21]).
@@
In Fig. 1 it is demonstrated that the Lieberstein model has solutions which are characteristic of an AP in earlier studies (for example, [1, 23]), however, it is known that in physiology, an AP has a few more properties that must be fulfilled. These are: (1) electrical nerve signal must annihilate upon head-on collision, (2) a disturbance or signal generated during a short time after the signal has passed a location (so-called refractory period) must dissipate rapidly without generating a new AP (a minimum separation time before another nerve pulse can propagate). In Figs. 2 and 3 it is demonstrated that the Lieberstein model satisfies these conditions. As we are using periodic boundary conditions (nodes and are connected so a wave hitting the boundary enters the grid again from , as an example). Figure 2 demonstrates that during a head-on collision of propagating AP signals annihilate according to the Lieberstein model fulfilling the first condition for a realistic AP model. The left panel of Fig. 2 depicts the AP amplitude in time at boundary node and the right panel depicts the spatial distribution of the signal when two spatial periods are plotted side by side (from to ) making it clearer to see a head-on collision of propagating AP signals at the right boundary. The left panel of Fig. 3 demonstrates the emergence of the AP from the initial disturbance (centred at ) above the threshold value given at and then later another voltage spike is given at the same location at which dissipates rapidly. The middle panel of Fig. 3 demonstrates that if a sufficient temporal separation between the pulses is given a repeating signal can be generated and the right panel of Fig. 3 shows a niche scenario where giving a small signal which is initially below the threshold value sufficiently frequently it is still possible to generate propagating AP if the combined dynamics succeeds in crossing the threshold value for the AP generation.
2.4 Influence of the axon radius on the propagation velocity of AP.
@@
From earlier studies (see, for example, [45]) it is known that the propagation velocity of the AP is influenced by three factors – (i) space coefficient , (ii) time coefficient , and (iii) the time measure characterising the time it takes to generate the AP on any point along the axon. The membrane resistance scales inverse-proportionally with axon diameter while membrane capacitance increases proportionally with axon diameter while axial resistance scales as meaning that the propagation velocity of AP is .
For a start, a so-called “sanity test" was carried out using the data of the HH celebrated experiment [1] In this case the temperature was 6.3 [Co] and we chose [mHcm]. Then the simulation yields an AP propagation velocity of 19.0 [m/s] compared to 18.8 [m/s] in the HH paper [1] against the experimentally measured AP velocity of 21.2 [m/s].
@@
We note that in the model parameters, the axon resistance is changed when the axon radius is changing so , where is constant [1]. The propagation velocity of AP as a function of axon radius is depicted in Fig. 4 with the parameters used for a numerical example. It should be noted that the electrical circuit hypothesis for the AP propagation velocity (like for the HH equations [1] or for the Lieberstein model [23]) which is based on similarity with the chosen electrical circuit, while certainly useful, can overlook some potentially important mechanisms (like, for example, the influence of cytoskeleton) that are left aside during the many simplifications needed to get a simple electrical scheme. This means that in practical applications one has to consider carefully any conflicts between experimental observations and modelling results to determine if the model is sufficient for adequately describing the observed process. For example, here we have chosen so that at [µ] the AP velocity is as a demonstration of the model behaviour, for practical applications that parameter would need to be determined from the electrophysiology of the problem under investigation as in reality, this is not a free parameter.
| Table 2: AP velocity as a function of axon radius . | ||||||||
| a µ | 0.25 | 0.50 | 1 | 2 | 4 | 8 | 16 | 25 |
| 0.5 | 0.75 | 1.0 | 1.4 | 1.95 | 2.55 | 3.2 | 3.55 | |
| a µ | 40 | 50 | 60 | 80 | 100 | 125 | 150 | 175 |
| 3.9 | 4.0 | 4.1 | 4.25 | 4.3 | 4.35 | 4.4 | 4.4 | |
@@ 3.1 The structure of myelinated axon
Nodes of Ranvier contain much higher densities of various ion channels than elsewhere on axons. K and Ca channels are about 10 nm of the radial length across the membrane and roughly 4 nm of the longitudinal diameter in the plane of the membrane [52]. Na channel is about 12 nm of the radial length across the membrane and about 10 nm of the longitudinal diameter in the plane of the membrane [53]. Structurally the myelinated part of the axon is divided into the following regions: next to the node of Ranvier is a region called the paranode. This is the area where the myelin attaches to the axon. The juxtaparanode is located next to the paranode and it is the area where most voltage-gated K+ ion channels are located. The Na+ channels are concentrated in the nodes of Ranvier. The geometry of a myelinated axon is sketched later in Fig. 5. A simplified model of the sheath can be presented as a stack of lipid bilayers. It should be noted that in the proposed model we do not consider different ion channel distributions between juxtaparanode and paranode and take uniform ion channel distribution within the node of Ranvier. The AP propagates down an axon without diminution at velocities up to 1 m/s (without myelin sheath) [29]. In non-myelinated neurons, the conduction velocity of an action potential is roughly proportional to the diameter of the axon [29]. The presence of a myelin sheath around an axon increases the velocity of impulse conduction to 10-100 m/s [29, 30]. This is a brief overview of the structure of the myelinated axons needed for further modelling. The detailed overview of the properties of myelinated nerve fibres is presented in [34, 54].

### v3

@@ 5 Solutions of Lieberstein model
reducing, therefore, the PDE to an ODE and then using standard ODE solvers for integration in time. For integration in time, the model equations are rewritten as a system of first-order ODEs and a standard numerical integrator is applied. In the numerical examples given in the present paper the ODEPACK FORTRAN code (see [22]) ODE solver is used through its NumPy implementation. Handling of the data and initialization of the variables is done in Python by making use of the package SciPy (see [27]) and the numerical results are analysed and visualised in the Matlab environment.
The following parameters are used: (number of spatial nodes), µ (membrane capacitance), µ (axon radius), (axoplasm resistance) while the HH model parameters are the same as taken in [23]: and while (initial values for parameters in HH equations at ). In the following example we take the axoplasm capacitance as µ and (resistance of an axon per unit length).
5.1 Estimating the value of parameter
We take (inductivity) for the numerical example (resulting in 21.2 propagation velocity for the AP signal with the same parameters as in the [23] which was experimentally observed value in that paper). For calculating the velocity of the AP from the numerical simulation we take a simple between maximum of the half-space at 5 [ms] and 10 [ms] (see Fig. 2) for profiles that propagate at less than 25 [m/s] and at 3 [ms] and 5 [ms] for profiles propagating faster than 25 [m/s]. Figure 2 is with parameters from HH classical paper [23] demonstrating the case where choosing [mHcm] yields the velocity for the AP [m/s] which was, as noted, the experimentally measured value. It should be noted that the classical HH model was giving AP velocity of 18.8 [m/s]. We remark that, as noted earlier by Kaplan and Trujillo [29], have calculated axon inductance from the movement of ions as if AP velocity is at µ but later in the same paper they also estimate a much smaller value for based on different physical considerations.
@@ 5.1 Estimating the value of parameter
It can be noted that the chosen value of inductance is large (22.2 [mHcm]) compared to, for example, a solid copper wire which would have self inductance of roughly 9 if the radius of the wire is 238 µ and the length is 1 . Although, clearly, the structure of the axon is very different from a solid metal fibre, the difference of roughly 6 orders of magnitude is significant. What inductivity means, in principle, can be described as a process that resists the change of current and considering various geometrical structures (like cytoskeleton) and different physical processes interacting with each other the chosen value does not appear to be outright implausible. One can note, for example, the significant mass difference between electrons and considered ions. However, it is clear that the nature of inductivity in the context of nerve fibres needs further clarification and study. We have a process in the proposed model that behaves like inductivity, and has a dimension related to inductivity so we have to make an assumption that it is inductivity to avoid violating Occam’s razor principle – although the relatively large value of the parameter needed to explain the experimental observations hints that there might be something more going on (see also Wang et al. [57]).
5.2 Numerical example
Initially we generate a narrow bell-shaped pulse (“spark”) for the in the middle of the space domain which generates the propagating AP. In spatial units, the length of the computation node () is 92 while the width of the computational domain in the space is 24. For the sake of readability, the additional equations and values of physical quantities used in numerical simulation are collected in Table 1 following [23] and [31]:
| Table 1: Parameters for numerical example (collected from [23] and [31]). | ||
It should be stressed, that the chosen parameters do not represent any particular model nerve as these have been collected from studies describing different experimental setups or even taken as a rough estimate (like specific inductivity ). Most of the parameters are taken from [23] corresponding to squid giant axon at about while inductivity is chosen so that the AP signal would have propagation velocity of around 21.1 if the axon radius is 238 [µ.
@@ 5.2 Numerical example
In Fig. 2 it is demonstrated that the Lieberstein model has solutions which are characteristic of an AP in earlier studies (for example, [23, 31]), however, it is known that in physiology, an AP has three key properties that must be fulfilled by any valid model aiming to descibe AP. These are: (1) there exist all-or-nothing threshold above which the AP is generated and under which the signal dissipates rapidly (Fig. 3), (2) electrical nerve signal must annihilate upon head-on collision (Fig. 4), (3) a disturbance or signal generated during a short time after the signal has passed a location (so-called refractory period) must dissipate rapidly without generating a new AP (a minimum separation time before another nerve pulse can propagate) (Fig. 5). In Figs. 3, 4 and 5 it is demonstrated that the Lieberstein model satisfies these conditions. Figure 3 demonstrates the existence of a threshold value for the Lieberstein model using the parameters mostly from [23] where initial excitation for the AP with amplitude of 11 [mV] dissipates rapidly while initial excitation with amplitude of 12 [mV] crosses the threshold with the used model parameters and generates the propagating AP signal. Figure 4 demonstrates that during a head-on collision the propagating AP signals annihilate according to the Lieberstein model fulfilling the second condition for a realistic AP model. The left panel of Fig. 4 depicts the AP profile in space with 5 [ms] intervals including the head-on collision of the profiles at 10 [ms] with the generation of two AP’s at and while the behaviour of the signals in space-time is easier to understand from the time-slice plot on the right panel where the evolution is visualised with 0.5 [ms] intervals with horisontal axis representing space and vertical axis time. Figure 5 demonstrates existence of refraction period for the considered model where a disdurbance with 20 [mV] amplitude (threshold value was approx 12 [mV]) fails to generate a propagating AP signal when given 12 [ms] after the initial pulse but suceeds generating the propagating AP signal when given 13 [ms] after the initial pulse. These examples together with the results of Lieberstein [31] complete the full analysis of an improved model for an AP in unmyelinated axon.
5.3 Influence of the axon radius on the propagation velocity of the AP.
@@
From earlier studies (see, for example, [35]) it is known that the propagation velocity of the AP is influenced by three factors – (i) space constant , (ii) time constant , and (iii) the time measure characterising the time it takes to generate the AP on any point along the axon. The membrane resistance scales inverse-proportionally with axon diameter while membrane capacitance increases proportionally with axon diameter while axial resistance scales as meaning that the propagation velocity of AP is .
@@
We note that in the model parameters, the axon resistance is changed when the axon radius is changing so , where is constant [23]. The propagation velocity of AP as a function of axon radius is depicted in Fig. 6 with the parameters used for a numerical example. It should be noted that the electrical circuit hypothesis for the AP propagation velocity (like for the HH equations [23] or for the Lieberstein model [31]) which is based on similarity with the chosen electrical circuit, while certainly useful, can overlook some potentially important mechanisms (like, for example, the influence of cytoskeleton) that are left aside during the many simplifications needed to get a simple electrical scheme. This means that in practical applications one has to consider carefully any conflicts between experimental observations and modelling results to determine if the model is sufficient for adequately describing the observed process. It should be stressed that here the method for calculating the velocity (see Fig. 2) of the AP profile is different than was used in the [23] or [29] where the authors used an velocity estimate for the moving frame of reference while here we have used the propagation velocity measured from the numerical solution (including the nonlinear effects) by tracking the coordinate of the maximum of the AP profile at two different time moments and then taking to find the propagation velocity.
6 Discussion and summary
Starting from the elementary form of Maxwell equations (1), (2) and drawing inspiration from the classical HH paper [23] we use the total membrane current density (6) to construct the model equations for unmyelinated axon similar to the Lieberstein model [31]. While Lieberstein opted to go a step further by moving into a moving frame of reference we opted to stay with the form which is closer to the Maxwell equations for a transmission line eqs. (8), (9) and used the work of Lieberstein as a source of inspiration for handling the question of inductance in the context of signal propagation along the nerve axon. Using the parameters from the earlier experimental studies we investigated briefly the solutions of the noted model for the unmyelinated axon and demonstrated that the behaviour of the solutions is in the physiologically plausible range and the key characteristics of the nervous signalling are fulfilled. These are: (i) the annihilation of AP signals during a head-on collision, (ii) the existence of activation threshold, and (iii) the refraction period after signal passing. In the parameter range considered, we have calculated the AP signal propagation velocity (see Fig. 6) from 0.68 (at [µm]) up to about 30.74 (at [µm]) for the unmyelinated axon. The key difference between the classical HH model and the Lieberstein-inspired model used here is that the mechanism for signal propagation along the axon emerges like a wave as a consequence of opting to keep the inductivity . While, indeed, there exist variations of the classical HH model which support AP signal propagation where the equations are written in the form of PDE instead of the usual ODE form (which describes signal evolution in time at a fixed spatial point). In these equations, normally, the potential-gradient-type member responsible for propagating the signal along the axon is not as clearly defined from the viewpoint of physics (usually some kind of abstracted diffusion-type process is used). We remark, that this is essential for the purpose of clearer physical interpretation as we make use of the model based on the elementary form of Maxwell equations which will be modified in the further studies to include the influence of myelination on the signal propagating along the axon.
Having a relatively simple pair of PDEs which are connected to the Maxwell equations for anything involving the movement of charges in an environment could be considered superior to investigating causal connections and making predictions than something that is not as clearly connected to the basic physical considerations.
To sum up, we revisited a model proposed by Lieberstein [31], solved a version of the model numerically with a physiologically viable set of parameters and checked that the solutions behave as an observed AP should. The results confirm the earlier studies [31, 29] with more details (like, for example, profiles of ion currents, and phenomenological variables). This forms a solid foundation for modifying the presented model for AP propagation in the future to include the effect of myelination (additional structure attached to the axon surface). The modification is in progress.
Acknowledgements
This research was supported by the Estonian Research Council (PRG 1227). Jüri Engelbrecht acknowledges the support from the Estonian Academy of Sciences.
Author contributions
Kert Tamm contributed by writing manuscript text and composing figures, performing numerical simulations and analysis of the results. Tanel Peets contributed analysis for comparing the modified Lieberstein equation against the classical Hodgkin-Huxley equations. All other work (like novel ideas presented, physical interpretation of the results, proofreading, etc) is a collaborative effort equally contributed by all authors.
Part 2 – The modelling of the action potentials in myelinated nerve fibres
7 Introduction
The celebrated Hodgkin-Huxley (HH) model can describe the action potential (AP) in an unmyelinated axon taking into account sodium and potassium ion currents [23]. The strength of the HH equations is in the detailed description of the physics of ion currents but the cable equation describing the propagation of an AP is simplified by neglecting the inductance. There are several ways to improve this classical model towards a better description of the structural properties of axons and taking into account the accompanying effects. Many studies are devoted to the description of mechanical and/or thermal effects accompanying the propagation of an AP [11, 9, 4, 28] resulting in the formation of an ensemble of waves. These theoretical models are based on experimental results [26, 51, 54, 59, 55, 52]. Another important avenue of studies is related to modelling the behaviour of an AP in myelinated axons. The starting point for the modelling of an unmyelinated axon is a cylindrical tube embedded into the extracellular fluid. The barrier between extra- and intracellular fluids is a lipid bilayer with proteins. Such a biomembrane is sometimes modelled like a simple lipid bi-layer only [11]. In the case of the myelinated axon, this barrier has a myelin sheath which consists of multiple layers of a glial membrane composed of lipids and proteins and serves as an insulator [56, 7] or in some scenarios even as signal modulating element [14]. It means that under the myelin sheath the ion currents through the basic biomembrane (barrier) proposed by Hodgkin and Huxley [23], are suppressed. However, the myelin sheath is interrupted by Ranvier nodes where the usual HH model works. It is proposed that under the myelin sheath, the passive cable equation (the diffusion-type equation) describes the process and in Ranvier nodes, the usual HH model can be applied [15, 20]. However, it is demonstrated experimentally that neurons which have myelin sheathing often show higher AP velocities compared to similar unmyelinated neurons. This effect was already reported by Lillie in 1924 [32] and is nowadays referred to as the Lillie transition [60] or saltatory conduction (see [25]).
In this paper, a phenomenological model is proposed for describing the propagation of an AP in a myelinated axon. The most important question is how to model the possible changes in the propagation velocity of an AP due to the myelin sheath. The process is nonlinear and we follow a recommendation of Whitham [58]: “… one should not always turn too quickly to a search for the .” In the context of our equation, it means that we should keep the neglected inductance as small as it is. This idea is supported by Wang et al [57] who have argued that inductance is “a missing piece of neuroscience”. They propose that the main source of the inductance is the myelin sheath. In general terms, however, they state for the inductance that “there is a kind of biological structure that can store energy in a non-electrical form” [57]. If we keep the inductance in the cable equation then, mathematically, under the myelin sheath the propagation is described not by a diffusion-type (see [20]) equation but by a hyperbolic equation. This hypothesis may better explain the changes in the velocity of an AP. Given reports [42] noting the dependence of velocity on the diameter of a fibre, this relationship should be taken into account.
Section 2 is introducing the model for the unmyelinated axon [31, 49]. Section 3 describes experimental results on myelinated axons focused on velocity changes [33, 44, 1], This analysis permits us to propose in Section 4 a phenomenological model describing the propagation of an AP in a myelinated axon influenced strongly by the structure of a myelin sheath (-ratio). The structure of the governing equations follows the Lieberstein model (see Section 2) where the inductance leads to a wave-like behaviour but the final velocity of an AP depends on ion currents. The numerical simulations demonstrate clearly the changes in the velocity under the myelin sheath of the propagating AP. Finally, in Section 5 the discussion is presented with conclusions. As far as the numerical simulations are carried on with physical variables (i.e., in physiologically viable range), the results could be better checked in experimental studies.
8 The modelling of an unmyelinated axon
We use a model describing AP propagation in unmyelinated axon (see Fig. 7) as a starting point [49, 31]. First, a governing equation for describing the AP
and, second governing equation for describing the current along the axon
Here is space (along the axon), is time, is the action potential, is the line axon current (along the axon), is the membrane current per unit length (expressed here through dimensionless parameters , channel conductances and equilibrium potentials [23, 31]), is the radius of the axon, is the axon resistance per unit length, is the axon specific self-inductance, is the axon self capacitance per unit area per unit length (often neglected as significantly smaller than but included here initially for the sake of completeness), is the membrane capacity per unit area. As is small then and the term can be neglected [31]. Equations (13) and (14) can be considered as a variant of the classical Hodgkin-Huxley model where the influence of specific inductance has not been neglected leading to a hyperbolic PDE (i.e., signal propagation velocity is finite) as opposed to parabolic (diffusion-type) governing equations as they arise in the HH model when inductance is dismissed. Equations (13) and (14) are taken as a starting point for modification to include the influence of myelination on the AP propagation dynamics along the axon.
It should be noted that the estimated value of needed to explain the experimental observations can be relatively large, at around 4420 [31, 36] as noted by Kaplan and Trujillo [29] if AP velocity is at µ. Cole and Baker have also investigated question of inductance in this context [5, 6]. It can be noted that similar models involving inductivity are in use in cardiophysiology, see, for example, [41]. Kaplan and Trujillo [29] have also derived inductance estimate from the inertia of the ions (note that Na and K ions are several orders of magnitude more massive than electrons, as carriers of charge) finding that estimate much smaller and conclude that if there is indeed that much inductance present it can not be explained purely by the mass of the carriers of charge and there must be some other effects involved as well. We remark that inductivity is, in principle, related to a process that resists too rapid change of the current and considering various geometrical structures (like cytoskeleton) and different physical processes interacting with each other the chosen value does not appear to be outright implausible [36]. However, it is clear that the exact nature of inductivity in the context of nerve fibres needs further clarification and study. We have a process in the proposed model that behaves like inductivity, and has a dimension related to inductivity so we have to make an assumption that it is inductivity to avoid violating Occam’s razor principle – although the relatively large value of the parameter needed to explain the experimental observations hints that there might be something more going on (see also Wang et al. [57]). We have opted to use for the numerical example (chosen to get 21.2 propagation velocity for the AP signal if axon radius is 238 [µ) and the rest of the parameters are like they were in the Hodgkin and Huxley classical paper where the HH model was introduced [23]. The 21.2 velocity was the experimentally observed propagation velocity and the original HH model provided velocity estimate of 18.8 [23].
It should be stressed that here the method for calculating the velocity (see Fig. 8) of the AP profile is different than was used in the [23] or [29] where the authors used an velocity estimate for the moving frame of reference while here we have used the propagation velocity measured from the solution (including the nonlinear effects) of the numerical simulation by tracking the coordinate of the maximum of the AP profile at two different time moments and then taking to find the propagation velocity.
@@ 9.1 The structure of myelinated axon
Nodes of Ranvier contain much higher densities of various ion channels than elsewhere on axons. K and Ca channels are about 10 nm of the radial length across the membrane and roughly 4 nm of the longitudinal diameter in the plane of the membrane [8]. Na channel is about 12 nm of the radial length across the membrane and about 10 nm of the longitudinal diameter in the plane of the membrane [46]. Structurally the myelinated part of the axon is divided into the following regions: next to the node of Ranvier is a region called the paranode. This is the area where the myelin attaches to the axon. The juxtaparanode is located next to the paranode and it is the area where most voltage-gated K+ ion channels are located. The Na+ channels are concentrated in the nodes of Ranvier. The geometry of a myelinated axon is sketched later in Fig. 9. A simplified model of the sheath can be presented as a stack of lipid bilayers. It should be noted that in the proposed model we do not consider different ion channel distributions between juxtaparanode and paranode and take uniform ion channel distribution within the node of Ranvier. As a rough ballpark AP propagation velocity for mammals could be taken at around 1 [m/s] (without myelin sheath) [33], however, for example, in the classical Hodgkin-Huxley paper where the HH model was initially introduced [23] the observed AP propagation velocity is 21.2 [m/s] for the giant axon of the squid. In non-myelinated neurons, the conduction velocity of an action potential is roughly proportional to the diameter of the axon [33]. The presence of a myelin sheath around an axon typically increases the velocity of impulse conduction to 10-100 [m/s] [33, 44]. This is a brief overview of the structure of the myelinated axons needed for further modelling. The detailed overview of the properties of myelinated nerve fibres can be found, for example, in [38, 40].